# Sparse Extended Mean-Variance-CVaR Portfolios with Short-selling

### v2

@@ 1 Introduction
The mean-variance CVaR problem is crucial in finance, as it provides a robust framework for optimizing portfolios by considering both expected returns (mean) and associated risks (variance and conditional value at risk). This approach helps investors balance the trade-off between maximizing returns and minimizing losses, enhancing portfolio performance and risk management in volatile markets [5, 13, 18]. Including short-selling is essential, as it allows investors to profit from downward market movements, improving portfolio diversification and risk management by providing additional opportunities for profit generation in declining markets [20].
Moreover, sparse optimization plays a critical role in various fields, such as machine learning, signal processing, and finance, offering efficient solutions for problems where the underlying data or parameters exhibit sparsity [3, 6, 19, 22, 21]. In portfolio optimization, having only a few assets can significantly impact diversification, potentially increasing exposure to individual asset risk and volatility. This highlights the importance of careful asset selection and risk management strategies. Hence, the cardinality constraints have been imposed in various portfolio optimization models [4, 11, 14].
@@
where $\lambda_{1}$ and $\lambda_{2}\geq 0$, and $\lambda_{3}=1-\lambda_{1}-\lambda_{2}\geq 0$ are the weights of variance, return, and CVaR, respectively. $\beta\in(0,1)$ is the confidence level, $\delta\geq 0$ is the cost per change in the proportion of the $i$-th asset, $L,U\in\mathbb{R}^{n}$ and $L<U$ are the lower and upper bounds for the weight vector of assets $x$, and $A\succeq 0$ is the covariance matrix. Also we have $\Psi=\mu+r_{c}h,\eta=\mu$, and $d_{j}:=b_{j}-r_{c}h;\ \forall j\in[m]$, where $\mu$ is the return vector, $r_{c}\geq 0$ is the risk-neutral interest, $h_{j}$ is the portion of risk free return for the investor when stock $i$ is on a short selling position. Finally, $b_{j},\ j\in[m]$ are the generated scenarios and $\phi_{i}$ is the proportion of the initial wealth invested in the $i$th stock.
Problem ($P$) has been the subject of several research [9, 16, 17]. When $\lambda_{1}=0$, it reduces to the widely studied mean-CVaR with a cardinality constraint, see, for example, [7, 10] and references therein. It is worth noting that by minimizing both Variance and CVaR simultaneously, we are essentially seeking a portfolio that exhibits low overall volatility (variance) and resilience to crises (CVaR). In the event of a major market shock (such as the 2020 COVID crash), the portfolio is explicitly designed to limit the average severity of losses. This is why we use two measures of risk together.
Recently, Hamdi et al. [9] applied the penalty decomposition method (PDM) to solve ($P$) and reported promising numerical results. Nevertheless, the full structure of the problem was not exploited. Therefore, we introduce a novel penalty decomposition algorithm customized to fully leverage every component of the objective function in ($P$) along with its constraints. We solve a sequence of penalty subproblems, of which a saddle point is efficiently identified using a BCD method. We meticulously examine each corresponding subproblem within this BCD method to derive closed-form solutions whenever feasible. Specifically, to handle the $\|.\|_{1}$ norm in the objective function along with bound constraints, we introduce a new variable. Additionally, another variable is required to manage the softmax function in the objective function. Lastly, we resort to a third variable to address the cardinality constraint, along with the short-selling constraints. Our discussions show how introducing these new variables leads to structured subproblems, which, in turn, admit closed-form or otherwise efficient solutions. We rigorously establish the convergence of our introduced algorithm toward a Lu–Zhang [12] minimizer of the non-differentiable problem ($P$) with non-convex constraints.
@@
Notation. The complement of a set $S$ is denoted as $S^{c}$. We use $|S|$ to represent its cardinality. For a natural number $n$, we define $[n]$ as the set $\{1,2,\dots,n\}$. Now, consider a set $S$ given by $S=\{i_{1},i_{2},\dots,i_{|S|}\}$, which is a subset of $[n]$. For any vector $x$ in $\mathbb{R}^{n}$, we denote the coordinate projection of $x$ with respect to the indices in $S$ as $x_{S};0]$, which means that the $i$th element of this vector equals $x_{i}$ when $i$ belongs to $S$, and it equals 0 for $i$ in $S^{c}$. We determine whether a matrix $A$ is positive semidefinite or definite by the notations $A\succeq 0$ and $A\succ 0$, respectively. In this paper, we let ${\rm sgn}(a):=1$ for $a>0,{\rm sgn}(a):=-1$ for $a<0,$ and ${\rm sgn}(0):=0$. For $x$ and $y\in\mathbb{R}^{n}$, $x\circ y$ shows the Hadamard (element-wise) multiplication of $x$ and $y$. Recall that [2]
@@ 3.2 Analysis of Algorithm 2
Because all constraints in ($P$) except the sparsity constraint are affine, the linearized operator has full row span. Hence, condition (18) holds automatically at any feasible point $x^{*}$. Moreover, since $\mathcal{X}=\mathbb{R}^{n}$, we have $T_{\mathcal{X}}(x^{*})=\mathbb{R}^{n}$. Hence, Robinson’s constraint qualification for a Lu–Zhang minimizer $(x^{*},\gamma^{*})$ of ($P$) requires the existence of an index set $\mathcal{L}\subseteq\{1,\dots,n\}$ with $|\mathcal{L}|=k$ and $x^{*}_{\mathcal{L}^{c}}=0$, ensuring that the following condition is satisfied [12]:
@@
Under Theorem 3.34 of [15] and Theorem 2.3 of [12], by virtue of the linearity of constraints in ($P$), except for the cardinality constraint, we conclude that Algorithm 2 converges to a Lu–Zhang minimizer of this problem under the given assumptions. ∎
@@ 4 Numerical Results
In this section, we compare the performance of Algorithm 2 with the algorithm in [9]. The direct approach of solving ($P$) using CVX-Mosek is also reported for gap computation. We use the data of the S&P index for 2018-2021, with 120 stocks when $\lambda_{1}=\lambda_{2}={1}/{3}$, $\phi=0$, $U_{i}=-L_{i}=0.2;\forall i\in[n],\delta=0.002$, and $\beta=0.95$. Moreover, we apply the Geometric Brownian Motion (GBM) model in [1] to generate scenarios. In the PADM algorithm of [9], the initial penalty parameter is set to $1.2$, and it is updated by a factor of 3. The inner loop is stopped when $(x^{s,l},w^{s,l})-(x^{s,l-1},w^{s,l-1})||_{1}\leq 10^{-5}$. It terminates with a partial minimum if $x-w||_{1}\leq 10^{-5}$. In Algorithm 2, the initial penalty parameter $\rho$ is set to $1.2$, and it is updated with the factor $r=3$. The inner loop is stopped for $\epsilon_{I}=10^{-5}$, and the outer loop is terminated when the convergence criterion $\epsilon_{O}=10^{-5}$ is met. All computations are performed in MATLAB R2017a on a 2.50 GHz laptop with 4 GB of RAM, and CVX 2.2 ([8]) is used to solve the optimization models. The results comparing returns, risks, Sharpe ratios, CVaR values, CPU times, and gaps for different $m$ values are reported in Table 1. The gap in this table is ${|f-\overline{f}|}/{(|f|+1)}$, where $f$ is the return (risk, Sharpe ratio, CVaR) for the direct solution approach, and $\overline{f}$ is the return (risk, Sharpe ratio, CVaR) for the PADM algorithm or Algorithm 2.
These results show that both Algorithm 2 and PADM are significantly faster than the direct solution approach. Further, Algorithm 2 is about twice as fast as the PADM while having competitive gaps in returns, risks, CVaR, and Sharpe ratios for all $m$ values. These results for $m=3000$ are also depicted in Figure 1 for different numbers of stocks, demonstrating the competitiveness of Algorithm 2. These results confirm that Algorithm 2 is a better alternative to the direct solution approach than the PADM in [9].

### v3

@@ 1 Introduction
The mean-variance CVaR problem is crucial in finance, as it provides a robust framework for optimizing portfolios by considering both expected returns (mean) and associated risks (variance and conditional value at risk). This approach helps investors balance the trade-off between maximizing returns and minimizing losses, enhancing portfolio performance and risk management in volatile markets [undefd, undefl, undefp]. Including short-selling is essential, as it allows investors to profit from downward market movements, improving portfolio diversification and risk management by providing additional opportunities for profit generation in declining markets [undefr].
Moreover, sparse optimization plays a critical role in various fields, such as machine learning, signal processing, and finance, offering efficient solutions for problems where the underlying data or parameters exhibit sparsity [undefb, undefe, undefq, undeft, undefs]. In portfolio optimization, having only a few assets can significantly impact diversification, potentially increasing exposure to individual asset risk and volatility. This highlights the importance of careful asset selection and risk management strategies. Hence, the cardinality constraints have been imposed in various portfolio optimization models [undefc, undefj].
@@
where $\lambda_{1}$ and $\lambda_{2}\geq 0$, and $\lambda_{3}=1-\lambda_{1}-\lambda_{2}\geq 0$ are the weights of variance, return, and CVaR, respectively. $\beta\in(0,1)$ is the confidence level, $\delta\geq 0$ is the cost per change in the proportion of the $i$-th asset, $L,U\in\mathbb{R}^{n}$ and $L<U$ are the lower and upper bounds for the weight vector of assets $x$, and $A\succeq 0$ is the covariance matrix. Also we have $\Psi=\mu+r_{c}h,\eta=\mu$, and $d_{j}:=b_{j}-r_{c}h;\ \forall j\in[m]$, where $\mu$ is the return vector, $r_{c}\geq 0$ is the risk-neutral interest, $h_{j}$ is the portion of risk free return for the investor when stock $i$ is on a short selling position. The sign constraint on $x$ is essentially $x_{i}\eta_{i}\geq 0;\forall i\in[n]$, which means for each stock that is not in a short-selling position, the proportion of investment is positive. Finally, $b_{j},\ j\in[m]$ are the generated scenarios and $\phi_{i}$ is the proportion of the initial wealth invested in the $i$th stock.
Problem ($P$) has been the subject of several research [undefh, undefn, undefo]. When $\lambda_{1}=0$, it reduces to the widely studied mean-CVaR with a cardinality constraint, see, for example, [undeff, undefi] and references therein. It is worth noting that by minimizing both Variance and CVaR simultaneously, we are essentially seeking a portfolio that exhibits low overall volatility (variance) and resilience to crises (CVaR). In the event of a major market shock (such as the 2020 COVID crash), the portfolio is explicitly designed to limit the average severity of losses. This is why we use two measures of risk together.
Recently, Hamdi et al. [undefh] applied the penalty decomposition method (PDM) to solve ($P$) and reported promising numerical results. Nevertheless, the full structure of the problem was not exploited. Therefore, we introduce a novel penalty decomposition algorithm customized to fully leverage every component of the objective function in ($P$) along with its constraints. We solve a sequence of penalty subproblems, of which a saddle point is efficiently identified using a BCD method. We meticulously examine each corresponding subproblem within this BCD method to derive closed-form solutions whenever feasible. Specifically, to handle the $\|.\|_{1}$ norm in the objective function along with bound constraints, we introduce a new variable. Additionally, another variable is required to manage the softmax function in the objective function. Lastly, we resort to a third variable to address the cardinality constraint, along with the short-selling constraints. Our discussions show how introducing these new variables leads to structured subproblems, which, in turn, admit closed-form or otherwise efficient solutions. We rigorously establish the convergence of our introduced algorithm toward a Lu–Zhang [undefk] minimizer of the non-differentiable problem ($P$) with non-convex constraints.
@@
Notation. The complement of a set $S$ is denoted as $S^{c}$. We use $|S|$ to represent its cardinality. For a natural number $n$, we define $[n]$ as the set $\{1,2,\dots,n\}$. Now, consider a set $S$ given by $S=\{i_{1},i_{2},\dots,i_{|S|}\}$, which is a subset of $[n]$. For any vector $x$ in $\mathbb{R}^{n}$, we denote the coordinate projection of $x$ with respect to the indices in $S$ as $x_{S};0]$, which means that the $i$th element of this vector equals $x_{i}$ when $i$ belongs to $S$, and it equals 0 for $i$ in $S^{c}$. We determine whether a matrix $A$ is positive semidefinite or definite by the notations $A\succeq 0$ and $A\succ 0$, respectively. In this paper, we let ${\rm sgn}(a):=1$ for $a>0,{\rm sgn}(a):=-1$ for $a<0,$ and ${\rm sgn}(0):=0$. For $x$ and $y\in\mathbb{R}^{n}$, $x\circ y$ shows the Hadamard (element-wise) multiplication of $x$ and $y$. Recall that [undefa]
@@ 3.2 Analysis of Algorithm 2
Because all constraints in ($P$) except the sparsity constraint are affine, the linearized operator has full row span. Hence, condition (18) holds automatically at any feasible point $x^{*}$. Moreover, since $\mathcal{X}=\mathbb{R}^{n}$, we have $T_{\mathcal{X}}(x^{*})=\mathbb{R}^{n}$. Hence, Robinson’s constraint qualification for a Lu–Zhang minimizer $(x^{*},\gamma^{*})$ of ($P$) requires the existence of an index set $\mathcal{L}\subseteq\{1,\dots,n\}$ with $|\mathcal{L}|=k$ and $x^{*}_{\mathcal{L}^{c}}=0$, ensuring that the following condition is satisfied [undefk]:
@@
Under Theorem 3.34 of [undefm] and Theorem 2.3 of [undefk], by virtue of the linearity of constraints in ($P$), except for the cardinality constraint, we conclude that Algorithm 2 converges to a Lu–Zhang minimizer of this problem under the given assumptions. ∎
@@ 4 Numerical Results
In this section, we compare the performance of Algorithm 2 with the algorithm in [undefh]. The direct approach of solving ($P$) using CVX-Mosek is also reported for gap computation. We use the data of the S&P index for 2018-2021, with 120 stocks when $\lambda_{1}=\lambda_{2}={1}/{3}$, $\phi=0$, $U_{i}=-L_{i}=0.2;\forall i\in[n],\delta=0.002$, and $\beta=0.95$. Moreover, we apply the Geometric Brownian Motion (GBM) model in [undef] to generate scenarios. In the PADM algorithm of [undefh], the initial penalty parameter is set to $1.2$, and it is updated by a factor of 3. The inner loop is stopped when $(x^{s,l},w^{s,l})-(x^{s,l-1},w^{s,l-1})||_{1}\leq 10^{-5}$. It terminates with a partial minimum if $x-w||_{1}\leq 10^{-5}$. In Algorithm 2, the initial penalty parameter $\rho$ is set to $1.2$, and it is updated with the factor $r=3$. The inner loop is stopped for $\epsilon_{I}=10^{-5}$, and the outer loop is terminated when the convergence criterion $\epsilon_{O}=10^{-5}$ is met. All computations are performed in MATLAB R2017a on a 2.50 GHz laptop with 4 GB of RAM, and CVX 2.2 ([undefg]) is used to solve the optimization models. The results comparing returns, risks, Sharpe ratios, CVaR values, CPU times, and gaps for different $m$ values are reported in Table 1. The gap in this table is ${|f-\overline{f}|}/{(|f|+1)}$, where $f$ is the return (risk, Sharpe ratio, CVaR) for the direct solution approach, and $\overline{f}$ is the return (risk, Sharpe ratio, CVaR) for the PADM algorithm or Algorithm 2.
These results show that both Algorithm 2 and PADM are significantly faster than the direct solution approach. Further, Algorithm 2 is about twice as fast as the PADM while having competitive gaps in returns, risks, CVaR, and Sharpe ratios for all $m$ values. These results for $m=3000$ are also depicted in Figure 1 for different numbers of stocks, demonstrating the competitiveness of Algorithm 2. These results confirm that Algorithm 2 is a better alternative to the direct solution approach than the PADM in [undefh].